OpenStudy (anonymous):

Find the value of x so that the line passing through (10, 5) and (x, 9) has a slope of -2. A. -2 B. 8 C. -8 D. 2

OpenStudy (anonymous):

b

OpenStudy (anonymous):

(9-5)/(x-10) = -2 therefore x = 8

OpenStudy (anonymous):

9-5 / x-10 = -2 -2x+20 = 4 2x = 16 x = 8